OpenStudy (anonymous):

Matrix Multiplication math help!! I should get it by now but I have no clue!!

OpenStudy (anonymous):

Find -5A+4B \[A = \left[\begin{matrix}6 & 1 \\ -4 & -6\\ 7 & -7\end{matrix}\right]\] \[B = \left[\begin{matrix}-5 & -1 \\ -3 & -8\\ 6 & 8\end{matrix}\right]\]

OpenStudy (aravindg):

First find -5A and 4B

OpenStudy (anonymous):

How would I find that? like I don't understand how to get the answer

OpenStudy (aravindg):

If you multiply a matrix by K then all its elements are multiplied by K. Here K=-5 for first matrix and 4 for second matrix.

OpenStudy (anonymous):

So I multiply each number in the first by -5 and the second by 4

OpenStudy (aravindg):

Yes.

OpenStudy (anonymous):

Ok what would you do after you multiply them?

OpenStudy (anonymous):

you add them together correct?

OpenStudy (anonymous):

@AravindG ?

OpenStudy (kainui):

Yeah, just add the corresponding elements, and that's it. It's just like you'd expect it to be really.

OpenStudy (anonymous):

well I did the multiplication and then added them and none of the answers are the same as mine?

OpenStudy (kainui):

Show me what you think the matrix -5A looks like and also what you think 4B looks like.

OpenStudy (anonymous):

Ok so a is \[A = \left[\begin{matrix}-30 & -5 \\ 20 & 30\\ -35 & 35\end{matrix}\right]\] B is \[B = \left[\begin{matrix}25 & 5 \\ 15 & 40\\ -30 & -40\end{matrix}\right]\]

OpenStudy (kainui):

Here, i did 4B, I think you're doing something wrong.\[4B=4\left[\begin{matrix}-5 & -1 \\ -3 & -8\\ 6 & 8\end{matrix}\right]=\left[\begin{matrix}-20 & -4 \\ -12 & -32\\ 24 & 32\end{matrix}\right]\]

OpenStudy (kainui):

I think you're just multiplying B by -5 instead of 4, that's all.

OpenStudy (anonymous):

no for b it would be \[\left[\begin{matrix}-20 & -4 \\ -12 & -32\\ 24& 32\end{matrix}\right]\]

OpenStudy (kainui):

Now just add them up.

OpenStudy (anonymous):

Ha I feel dumb I should know this!! thank you guys so much!!

OpenStudy (kainui):

No problem! =)